### v2

@@ 3.2 Realising various 3d CFTs
The second direction is to study various other CFTs beyond 3d Ising. The fuzzy sphere has revealed lots of new information about these theories ; the previously known results are also consistent with the fuzzy sphere. So far, the accessible CFTs include $\mathrm{SO}(5)$ deconfined criticality [4], $\mathrm{O}(3)$ Wilson-Fisher [8], a series of new theories with $\mathrm{Sp}(N)$ symmetry [16], and the 3-state Potts model [18].
@@
The first theory besides Ising CFT to which fuzzy sphere is applied is the $\mathrm{SO}(5)$ deconfined quantum critical point (DQCP). Deconfined quantum critical point (DQCP) is one of the pioneering examples of phase transitions beyond Landau paradigm [62, 63, 64]. It has led to numerous theoretical surprises, including the emergent $\mathrm{SO}(5)$ symmetry [65] and the duality between interacting theories [66]. Despite extensive studies over the past two decades, its nature remains controversial. Numerical simulations have shown no signal of discontinuity, but abnormal scaling behaviours have been observed [64]. A plausible proposal to reconcile the tension is that DQCP is pseudocritical, i.e., a weakly first-order phase transition that has approximate critical behaviour, and is controlled by a pair of complex fixed points very close to the pseudocritical region [66, 67].
[…]
The $\mathrm{O}(3)$ Wilson-Fisher [8]
The $\mathrm{O}(N)$ Wilson-Fisher (WF) theories are probably one of the most studied theories for 3d criticalities with a wide range of applications [73, 26, 25]. Specifically, this work focuses on the $\mathrm{O}(3)$ WF CFT. The construction involves two copies of $\mathrm{SU}(2)$ ferromagnet with altogether four flavours. Briefly speaking, the model contains two competing terms : (1) a $\mathrm{SU}(2)$ ferromagnetic interaction which favours a Heisenberg ferromagnetic phase where each of the two copies is half-filled and the symmetry-breaking order parameter lives on a $S^{2}$ manifold, (2) a polarising term which favours one of the two copies being completely filled, corresponding to a Heisenberg paramagnet. The transition between these two phases falls into the $\mathrm{O}(3)$ Wilson-Fisher universality. Through the energy spectrum at the transition, Han et al. provide evidence that $\mathrm{O}(3)$ Wilson-Fisher fixed point exhibits conformal symmetry, as well as revealing a wealth of information about the CFT, e.g., the instability to cubic anisotropy. They also calculate several OPE coefficients.
@@
[…]
$\frac{\mathrm{Sp}(N)}{\mathrm{Sp}(M)\times\mathrm{Sp}(N-M)}.$ (5)
[…]
The 3-state Potts model [18]
The Potts models describe transitions that spontaneously break $S_{Q}$ symmetries where $Q\in\mathbb{Z}$ is known as the number of states. In 2d, the transitions with $Q\leq Q_{c}=4$ are continuous and captured by CFTs [75], while $Q>Q_{c}$ are first order. Specifically, 2d 5-state Potts transition is pseudocritical and described by a pair of complex CFTs in its vicinity [76] in a similar manner with the conjectured $\mathrm{SO}(5)$ DQCP. In 3d, the 3-state Potts model is found to be first-order [77, 78]. This work constructs a 3-flavour model on the fuzzy sphere with $S_{3}$ permutation symmetry among flavours. The interacting Hamiltonian resembles the Ising model. Interestingly, Yang et al. find out that the transition point of the 3d 3-state Potts model, despite being probably first-order, exhibits approximate conformal symmetry, indicating that there might be an underlying CFT describing it. However, it is difficult to determine the nature of the transition from the operator spectrum (specifically, from the relevance of the second singlet $\epsilon^{\prime}$) due to the complicated finite-size effect.
@@ 3.3 Studying conformal defects and boundaries
Apart from the bulk CFTs, the fuzzy sphere can also be used to study their conformal defects and boundaries. Deforming a CFT with interactions living on a sub-dimensional defect may trigger a RG flow towards a non-trivial interacting IR fixed point. A defect IR theory with a smaller conformal symmetry is called a defect CFT [79, 80]. The dCFTs have rich physical structures, such as defect operators and bulk-to-defect correlation functions. Moreover, a bulk CFT can flow to several different dCFTs. Similarly, deformation on the boundary may trigger a flow towards a boundary CFT (bCFT). So far, the accessible defects/boundaries include the magnetic line defect of the 3d Ising CFT, in particular, its defect operator spectrum, correlators [6], $g$-function, defect-changing operators [9], its cusp [11], and the conformal boundaries of the 3d Ising CFT [12, 13].
Conformal magnetic line defect [6]
This is the first work that studies conformal defects with the fuzzy sphere. The simplest example of conformal defect is the magnetic line defect of the 3d Ising CFT [81, 82, 83, 84], where the defect line is completely polarised and the $\mathbb{Z}_{2}$ symmetry is explicitly broken. A defect line along $z$-direction that passes the origin point, after the radial quantisation, corresponds to the north and south poles of the sphere being polarised. Hence, to realise the magnetic line defect on the fuzzy sphere, one only needs to apply a pinning magnetic field to the north and south poles (Since only the $m=+s$ orbital has non-zero amplitude at the north pole and $m=-s$ at the south pole due to the locality, one only need to pin the $m=\pm s$ orbitals).
@@
This work studies the $g$-function of conformal defects and the defect-creation and changing operators. Similar to the central charge and the $F$-function in bulk CFTs, there exists a RG-monotonic quantity called the $g$-function for the line defects that is non-increasing along the flow [85, 86]. It is defined as the ratio between the partition functions of the defect CFT and the bulk CFT. On a different note, consider two semi-infinite magnetic line defects pinned towards opposite directions joint at one point, a defect-changing operator lives at the joining point [87, 88]. Similarly, a defect-creation operator lives at the endpoint of a semi-infinite line defect. The relevance of the defect-changing operator is related to the stability of spontaneous symmetry breaking (SSB) on the line defect.
Zhou et al. realise the defect-creation and changing operators for the Ising magnetic line defect by acting a pinning field at the north pole, and opposite pinning fields at the north and south poles, respectively. The scaling dimensions are calculated through state-operator correspondence $\Delta_{\textrm{creation}}=0.108(5),\Delta_{\textrm{changing}}=0.84(5)$, indicating the instability of SSB on the Ising magnetic line. Moreover, they show that the $g$-function and many other CFT data can be calculated by taking the overlaps between the eigenstates of different defect configurations. Notably, this paper has given the first non-perturbative result for the $g$-function $g=0.602(2)$.
@@
A cusp is two semi-infinite defect lines joined at one point at an angle. This can be realised on the fuzzy sphere through pinning fields at two points at an angle. Cuomo et al. study the cusps through various theoretical and numerical approaches. In particular, on the fuzzy sphere, they calculate the cusp anomalous dimension as a function of the angle for the Ising magnetic line defect, and verify its relation with the Zamolodchikov norm of the displacement operator.
@@
3.4 Other works on the fuzzy sphere
Besides the three directions of works, several other works push the frontier of our knowledge of and techniques for the fuzzy sphere.
Conformal perturbation [5]
The energy spectrum calculated numerically at a finite size does not coincide with that of the CFT. Part of the finite-size correction comes from the higher irrelevant operators that are not precisely tuned to zero (e.g., in the Ising CFT, the irrelevant operators include $\epsilon^{\prime},C_{\mu\nu\rho\sigma},T^{\prime}_{\mu\nu}$, etc., and the lowest singlets $\epsilon$ and $\epsilon^{\prime}$ are tuned away through the two parameters). These irrelevant operators exert perturbations on the states and their energies. This paper captures this kind of correction using the conformal perturbation theory. By making use of the fact that the corrections from an irrelevant operator on the energy of the primary and its descendants are not independent, the coefficients of the irrelevant operators can be fitted and their corrections can be removed.
Although this work is not exactly a study on the fuzzy sphere, it opens up a new route to improving the precision of scaling dimensions on the fuzzy sphere by making better use of the existing data. The method to partly remove the finite-size correction through conformal perturbation theory is widely used by the following works.
Quantum Monte Carlo on fuzzy sphere [7]
Until this work, the numerical methods that have been applied to the fuzzy sphere include exact diagonalisation (ED) and density matrix renormalisation group (DMRG). Hofmann et al. further present the numerical studies of the fuzzy sphere with quantum Monte Carlo (QMC) simulation, which is known for its potential for studying criticalities in $(2+1)$ dimensions at larger system size. Specifically, they make use of the determinant quantum Monte Carlo (DQMC) method that converts the simulation of fermions into the simulation of bosonic auxiliary fields. To overcome the sign problem, they consider two copies of the original model and construct the Ising CFT on a 4-flavour model. They determine the lowest energy spectra within each symmetry sector by calculating the time-displaced correlation functions. They also calculate the equal-time correlation functions and compare them with the 2-pt functions of CFT.
Ising CFT from FQHE state [17]
Until this work, all the constructions of CFTs on the fuzzy sphere were based on the quantum Hall ferromagnet. Specifically, in the absence of the interaction, an integer number of the lowest Landau levels are fully occupied. This state has a finite charge gap that guarantees that the gapless spin degree of freedom does not strongly couple with the charge degree of freedom when one adds the interactions.
Voinea et al. further explore the possibility of constructing CFTs on other states with charge gap — in particular, the Haldane-Laughlin states that capture the fractional quantum Hall effect (FQHE) [47, 92]. Specifically, they study the fermionic LLL at fillings of $\nu=1/3$ and $1/5$. The model Hamiltonian contains (1) a dominant projection term that puts the ground state on the Haldane-Laughlin state, and (2) an interaction term as a perturbation that drives the Ising-type phase transition. They show that the energy spectra at the critical point exhibit conformal symmetry. Notably, they also make the construction with respect to the bosonic LLL at a filling of $\nu=1/2$.
@@ 3.4 Other works on the fuzzy sphere
Besides the fuzzy sphere $S^{2}$, the regularisation with the lowest Landau level can also be used for other manifolds or dimensions. This work studies the 2d CFT on a ‘fuzzy circle’. Although Landau levels can only be defined on even space dimensions [93], one can reach odd space dimensions by compactifying one of the even dimensions. Specifically, Han et al. construct the LLL on a thin torus $T^{2}$ and sends one of the lengths to infinity while fixing the other. In this way, a circle $S^{1}$ is recovered in the thermodynamic limit. They construct the 2d Ising and 3-state Potts CFTs on the fuzzy circle and compare the operator spectrum and OPE coefficients with the Virasoro multiplet structure and the known results in the exactly solvable minimal models.

### v3

@@ 3.3 Realising Various 3D CFTs
The third direction is to study various other CFTs beyond 3D Ising. The fuzzy sphere has revealed lots of new information about these theories ; the previously known results are also consistent with the fuzzy sphere. So far, the accessible CFTs include the 3-state Potts model [18], the Yang-Lee non-unitary CFT [21, 22, 23] and three widely studied classes :
  1. 1.
    the free-scalar [24, 25] and Wilson-Fisher CFTs realised as Heisenberg bilinear [8] and truncated quantum rotor model [32, 33],
  2. 2.
    CFTs with $\mathrm{Sp}(N)$ global symmetry related to the non-linear $\sigma$ models (NLSM) with a Wess-Zumino-Witten (WZW) topological term, including $\mathrm{SO}(5)$ deconfined criticality [4] and $\mathrm{O}(4)$ deconfined criticality [26] through a symmetry-breaking perturbation, a series of new theories with $\mathrm{Sp}(N)$ symmetry [16], and the multi-flavour $\mathrm{SU}(2)$ QCD [34].
While the CFTs above are all built from quantum Hall ferromagnets at integer filling, another class is
  1. 3.
    fractional quantum Hall transitions, which will be discussed in more detail in Section 3.4.
The 3-state Potts model [18]
The Potts models describe transitions that spontaneously break $S_{Q}$ symmetries where $Q\in\mathbb{Z}$ is known as the number of states. In 2D, the transitions with $Q\leq Q_{c}=4$ are continuous and captured by CFTs [80], while $Q>Q_{c}$ are first order. Specifically, 2D 5-state Potts transition is pseudo-critical and described by a pair of complex CFTs in its vicinity [81] in a similar manner with the conjectured $\mathrm{SO}(5)$ DQCP. In 3D, the 3-state Potts model is found to be first-order [82, 83]. This work constructs a 3-flavour model on the fuzzy sphere with $S_{3}$ permutation symmetry among flavours. The interacting Hamiltonian resembles the Ising model. Yang et al. find out that the transition point of the 3D 3-state Potts model, despite being probably first-order, exhibits approximate conformal symmetry, indicating that there might be an underlying CFT describing it. However, it is difficult to determine the nature of the transition from the operator spectrum (specifically, from the relevance of the second singlet $\epsilon^{\prime}$) due to the complicated finite-size effect.
Yang-Lee non-unitary CFT [21, 22, 23]
The Yang-Lee singularity [84, 85, 86] is one of the simplest non-unitary CFTs. It is a natural extension of the Ising CFT triggered by the $i\sigma$ deformation, and plays an important role in understanding order-disorder transformation. While the 2D Yang-Lee theory is described by $M(2,5)$ minimal model and 3D Yang-Lee CFT can be solved perturbatively up to five loops, Fan et al., Arguello Cruz et al., and Elias Miro et al. present the non-perturbative results of the Yang-Lee CFT on the fuzzy sphere. They have studied the operator spectrum, the OPE coefficients, and the RG flow from the Ising CFT to the Yang-Lee CFT. They have also shown how the finite-size effects can be controlled by finite-size scaling and conformal perturbation.
3.3.1 Free-Scalar and Wilson-Fisher CFTs
The $\mathrm{O}(N)$ vector model is arguably the most well studied theory that describes the critical behaviours. It can be formulated as a quantum field theory (QFT) of $N$ interacting scalar fields. This QFT has an ultra-violet fixed point of $N$ free scalars. A quartic interaction term induces a RG flow to the infra-red fixed point described by the Wilson-Fisher (WF) CFT [87, 41, 40]. It describes a transition between an ordered phase with $\mathrm{O}(N)$ spontaneous symmetry breaking and an $\mathrm{O}(N)$-symmetric disordered phase. This theory has been extensively studied using perturbative calculations, quantum Monte Carlo and conformal bootstrap. The $\mathrm{O}(2)$ WF has been experimentally measured with high precision in the normal-superfluid transition of Helium-4 [88]. Its fuzzy-sphere realisation paves the way for studying operators with certain quantum numbers, multi-scalar CFTs, defects and boundaries, etc.
The free scalar CFT [24, 25]
This work introduces a simple model to realise the free real scalar CFT on the fuzzy sphere that is structurally similar to the 3D Ising CFT. A weakly-broken $\mathrm{U}(1)$ symmetry in the fuzzy-sphere model realises the $\mathbb{R}$-shift symmetry of the free scalar in the thermodynamic limit, so that the fix point can be accessed with only a single tuning parameter. He numerically demonstrates that the model correctly reproduces the operator spectrum, correlation functions and the harmonic oscillator algebra of the real-scalar CFT. He generalises the Girvin-MacDonald-Platzman algebra [89] to the fuzzy-sphere algebra of the density operators, which is potentially useful for defining quantum field theories on non-commutative geometries. He proposes a wave-function ansatz for the ground states which exhibit remarkable agreement with the CFT ground state wave-functions of the fuzzy-sphere model.
The bilayer Heisenberg transition [8]
This work studies the $\mathrm{O}(3)$ WF CFT on the set-up of a bilayer Heisenberg model. The construction involves two copies of $\mathrm{SU}(2)$ ferromagnet with altogether four flavours. Briefly speaking, the model contains two competing terms : (1) a $\mathrm{SU}(2)$ ferromagnetic interaction which favours a Heisenberg ferromagnetic phase where each of the two copies is half-filled and the symmetry-breaking order parameter lives on a $S^{2}$ manifold, (2) a polarising term which favours one of the two copies being completely filled, corresponding to a Heisenberg paramagnet. The transition between these two phases falls into the $\mathrm{O}(3)$ Wilson-Fisher universality. Through the energy spectrum at the transition, Han et al. provide evidence that $\mathrm{O}(3)$ Wilson-Fisher fixed point exhibits conformal symmetry, as well as revealing a wealth of information about the CFT, e. g. the instability to cubic anisotropy. They also calculate several OPE coefficients.
The general model for $\mathrm{O}(N)$ free-scalar and Wilson-Fisher CFTs [33]
This work proposes a general model that realises the $\mathrm{O}(N)$ free-scalar and Wilson-Fisher CFTs. The set-up contains altogether $(N+1)$ flavours with total filling $\nu=1$, where one flavour $c_{0}$ is $\mathrm{O}(N)$ singlet and the rest $N$ flavours $c_{i}$ transform as $\mathrm{O}(N)$ vector. The scalar field is realised as the bilinear $\phi_{i}=c_{i}^{\dagger}c_{0}+c_{0}^{\dagger}c_{i}$, and the Hamiltonian for the WF CFT contains (1) a particle density interaction, (2) a $\phi$ density interaction, and (3) a relative chemical potential of the singlet flavour. Tuning the chemical potential realises a phase transition between a paramagnetic phase with $c_{0}$ fully filled and an $\mathrm{O}(N)$ symmetry-breaking phase where electrons are in a superposed state between a vector flavour and the singlet flavour. Turning to the $\mathrm{O}(N)$ free-scalar theory, it corresponds to a pseudo-Goldstone phase of a $\mathrm{PO}(N)$ SSB.
Guo et al. provide numerical evidence for $N=2,3,4$ such as the operator spectra and two-point conformal correlators. They discuss several quantities that may not have a direct CFT interpretation, but imply that the CFTs become to some extent semi-classical once regularised on the fuzzy sphere.
The detailed study of the $\mathrm{O}(3)$ Wilson-Fisher CFT [32]
This work focuses on the $\mathrm{O}(3)$ WF and fomulate the four-flavour model as a truncated rotor model. Dey et al. study the model numerically using ED and DMRG. They locate the critical point through a careful conformal perturbation analysis. They obtain scaling dimensions from finite-size spectra and OPE coefficients through conformal perturbation. The results are benchmarked with conformal bootstrap and perturbative calculations.
3.3.2 NLSM-WZW with Symplectic Symmetry
One route to realise many CFTs is through the non-linear sigma model (NLSM) with a Wess-Zumino-Witten (WZW) term. The NLSM on a Grassmannian
$\frac{\mathrm{G}(N)}{\mathrm{G}(M)\times\mathrm{G}(N-M)},\qquad\mathrm{G}(N)=\mathrm{Sp}(\tfrac{N}{2}),\mathrm{O}(N),\mathrm{U}(N),\dots$
captures a symmetry breaking pattern of fully filling $M$ out of $N$ flavours of fermions with a global symmetry $\mathrm{G}(N)$. The NLSM often allows a topological WZW term with a quantised level. On the Landau level, a WZW term naturally appear [90, 34]. While the 3D NLSM in its renormalisable region can only flow to the SSB phase, the symmetry and anomaly of the NLSM-WZW can be matched with critical gauge theories with critical scalars or fermions coupled to a dynamic gauge field [91]. These gauge theories can either serve as the UV completion of the NLSM, or has the same UV completion with the NLSM. The simplest ones to construct on the fuzzy sphere are the NLSM on symplectic Grassmannians (i. e. with $\mathrm{G}(N)=\mathrm{Sp}(\tfrac{N}{2})$) with a level-1 WZW term. [^4] The construction for $\mathrm{U}$ or $\mathrm{O}$ faces the difficulty that their centres $\mathrm{U}(1)$ or $\mathbb{Z}_{2}$, as the electric charge conservation or its subgroup, must be decoupled ; the construction for $\mathrm{SU}$ or $\mathrm{SO}$ faces the difficulty in realising the topological $\mathrm{U}(1)$ or $\mathbb{Z}_{2}$ symmetry. They can be constructed by considering $N$ flavours of fermions and break the flavour symmetry to $\mathrm{Sp}(\tfrac{N}{2})$. This construction can be used to realise various 3D QCDs.
@@ 3.3.2 NLSM-WZW with Symplectic Symmetry
Deconfined quantum critical point (DQCP) is one of the pioneering examples of phase transitions beyond Landau paradigm [92, 93, 94]. It has led to numerous theoretical surprises, including the emergent $\mathrm{SO}(5)$ symmetry [95] and the duality between interacting theories [96]. Despite extensive studies over the past two decades, its nature remains controversial. Numerical simulations have shown no signal of discontinuity, but abnormal scaling behaviours have been observed [94]. A plausible proposal to reconcile the tension is that DQCP is pseudo-critical, i. e. a weakly first-order phase transition that has approximate critical behaviour, and is controlled by a pair of complex fixed points very close to the pseudo-critical region [96, 97].
The DQCP can be conveniently studied on the fuzzy sphere by constructing a NLSM on $S^{4}$ with WZW level-1, which serves as a dual description of the DQCP with an exact $\mathrm{SO}(5)$ symmetry [95, 96]. The idea is to construct a 4-flavour model with symmetry $\mathrm{Sp}(2)/\mathbb{Z}_{2}=\mathrm{SO}(5)$ ($\mathbb{Z}_{2}$ means to project the pseudo-real representations). At half-filling, it can be described by a NLSM on the Grassmannian $\tfrac{\mathrm{Sp}(2)}{\mathrm{Sp}(1)\times\mathrm{Sp}(1)}\cong S^{4}$ and the WZW level can be matched [98, 99]. This work provides evidence that the DQCP exhibits approximate conformal symmetry. Zhou et al. have identified 19 conformal primaries and their 82 descendants. Furthermore, by examining the RG flow of the lowest symmetry singlet, they demonstrate that the DQCP is more likely pseudo-critical, with the approximate conformal symmetry plausibly emerging from nearby complex fixed points. Several works [100, 101] appear later to follow up.
The $\mathrm{O}(4)$ deconfined criticality [26]
Apart from the $\mathrm{SO}(5)$ deconfined criticality, one can reach the $\mathrm{O}(4)$, or easy-plane DQCP by adding a perturbation in the $\mathrm{SO}(5)$ symmetric tensor representation that breaks the global symmetry to $\mathrm{O}(4)$. The $\mathrm{O}(4)$ DQCP appears in several lattice models and has several gauge theory descriptions, e. g. QED with two flavours of Dirac fermions [96]. This work studies the RG flow from the $\mathrm{SO}(5)$ to the $\mathrm{O}(4)$ DQCP and traces the $\mathrm{O}(4)$ operators back to the $\mathrm{SO}(5)$ theory.
@@
The quest to discover new 3D CFTs has been intriguing for physicists. A virgin land on this quest is the parity-breaking CFTs. In 3D, the Chern-Simons-matter theories stand out as the most well-known and possibly the only known type of parity-breaking CFTs. The fuzzy sphere is a promising platform for studying these theories. This work makes a concrete construction by generalising the DQCP to the WZW-NLSM on the target space of a general symplectic Grassmannian $\frac{\mathrm{Sp}(N)}{\mathrm{Sp}(M)\times\mathrm{Sp}(N-M)}$. Several candidate Chern-Simons-matter theories are known to exist on its phase diagram with $N$ flavour of gapless bosons or fermions coupled to a non-Abelian (viz. $\mathrm{Sp}(1)$, $\mathrm{Sp}(2)$, etc.) Chern-Simons gauge field [91]. On the fuzzy sphere, this WZW-NLSM can be realised by a $2N$ layer model with $\mathrm{Sp}(N)$ flavour symmetry, and $2M$ out of the $2N$ layers are filled. Zhou et al. numerically verify the emergent conformal symmetry by observing the integer-spaced conformal multiplets and studying the finite-size scaling of the conformality.
Multi-flavour $\mathrm{SU}(2)$ QCD [34]
The conformal window is an outstanding problem for 3D gauge theories. For a given gauge field, when the number of flavours coupled to it exceeds some critical value $N>N_{c}$, the field theory flows to a conformal fixed point ; otherwise the gauge field is confined and the field theory flows to SSB. The conformal windows of QED3 and $\mathrm{SU}(2)$ QCD3 are directly related to the nature of deconfined criticality and the Dirac spin liquid. However, results beyond perturbative calculations for certain gauge groups are still hard to access.
On the fuzzy sphere, the $\mathrm{SU}(2)$ QCD3 with $N$ flavours of fermions can be realised with the $\mathrm{Sp}(N)$ model at half-filling $M=N/2$. The model is free of sign problem and thus accessible through quantum Monte Carlo. Notably, $N$ is treated as a parameter instead of the number of fermions simulated, so the calculation can reach arbitrarily large number of flavours.
This work shows that the phase diagram at $N\geq 4$ has a stable conformal fixed point corresponding to the QCD3 through examining the RG flow. Huffmann et al. extracts the various scaling dimensions from the equal-time and imaginary-time two-point correlation functions and finds agreement with conformal symmetry and large-$N$ expansion result.
3.4 Exploring Fractional Quantum Hall Transitions
The models above are all built from quantum Hall ferromagnets at integer filling. Specifically, in the absence of the interaction, an integer number of lowest Landau levels are fully occupied. The lowest Landau level at fractional filling realises various topological orders (TO) with long-range entanglement and anyon excitations. E. g., the Laughlin states [102] with filling $\nu=1/k$ realise the Abelian topological orders captured by Chern-Simons theories $\mathrm{U}(1)_{-k}$, the Jain sequence [103] with $\nu=p/(mp+1)$ realises the Abelian topological orders with composite fermion descriptions, and the Read-Rezayi sequence with $\nu=p/(mp+2)$ [104], as a natural extension of the Moore-Read states [105], realises in general a series of non-Abelian topological orders. These states have finite charge gaps as well that guarantees that the gapless spin degree of freedom can decouple from the strongly non-commutative charge degree of freedom.
The fuzzy sphere provides a suitable platforms to study the transitions between or out of these fractional quantum Hall (FQH) states. Theoretically, these transitions are often described by Chern-Simons theories coupled to matter, many of which are conjectured to be conformal with enhanced symmetries and field theory dualities [106, 107, 108]. The Moiré materials [109, 110, 111] provide an experimental platform to realise these transitions. The critical points can be conveniently reached by tuning many knobs like potential amplitude.
The exploration starts at realising the Ising CFT on the background of a FQHE state [17], for which the topological order stays the same across the transition. The transition between $\nu=1/2$ bosonic Laughlin state and a $\nu=2$ fermionic Laughlin state, realising the confinement transition of $\nu=1/2$ bosonic Laughlin state [27], is the first fractional quantum Hall transition on the fuzzy sphere. Umsing similar strategy, the phase diagram between fermionic integer quantum Hall and bosonic Laughlin state realising two transitions of free Majorana fermion and gauged Ising CFT [29], and the transition between bosonic Pfaffian and Halperin 220 state [30], are also studied.
Ising CFT from FQHE state [17]
Voinea et al. explore the possibility of constructing CFTs on the Haldane-Laughlin states that capture the fractional quantum Hall effect (FQHE) [63, 112]. Specifically, they study the fermionic LLL at fillings of $\nu=1/3$ and $1/5$. The model Hamiltonian contains (1) a dominant projection term that puts the ground state on the Haldane-Laughlin state, and (2) an interaction term as a perturbation that drives the Ising-type phase transition. They show that the energy spectra at the critical point exhibit conformal symmetry. Notably, they also make the construction with respect to the bosonic LLL at a filling of $\nu=1/2$.
The confinement transition of $\nu=1/2$ bosonic Laughlin state [27]
This work studies one of the simplest instances of Chern-Simons-matter theories : one complex critical scalar coupled to a $\mathrm{U}(1)_{2}$ Chern-Simons gauge field and its three other dual Lagrangian descriptions [108, 113, 114]. They describe the transition between a $\nu=1/2$ bosonic Laughlin state and a trivially gapped phase. This transition appears in various contexts in condensed-matter physics, viz. Kalmeyer-Laughlin chiral spin liquid [115, 116], anyon superconductivity [117, 118, 119], Feshbach resonance [120, 121], etc.
Zhou et al. realise this theory on a set-up with fermion-boson mixture containing two flavours of fermion and one flavour of bosons carrying electric charge $Q_{b}=2Q_{f}$. Tuning the relative chemical potential induces a transition between a $\nu_{f}=2$ fermionic integer quantum Hall state and a $\nu_{b}=1/2$ bosonic fractional quantum Hall state. They show that the transition is continuous and has emergent conformal symmetry. The operator spectrum has exactly one relevant singlet with scaling dimension $\Delta_{S}=1.52(18)$, signature of a critical point. This work opens up the route of using the fuzzy sphere to study transitions between distinct topological order.
The free Majorana fermion theory [29]
The CFTs realised on the fuzzy sphere until this work contains only bosonic CFT operators. The microscopic fermions on the fuzzy sphere have strong non-commutativity, transform under projective representations of the sphere rotation symmetry, and thus cannot flow to CFT operators. This work overcomes this challenge by constructing a boson-fermion mixed set-up with both microscopic bosons $b$ and fermions $f$ with the same electric charge and let their angular momenta differ by $1/2$. Their bilinears $b^{\dagger}f$ or $f^{\dagger}b$ can be used to realise fermionic local operators in a CFTs.
Zhou et al. allow conversion between two bosons and two fermions and set the total filling $\nu=1$. On the phase diagram, there are three gapped phases, viz. a fermionic integer quantum Hall phase, an $f$-wave chiral topological superconductor, and a bosonic Pfaffian phase [105]. They are separated by two continuous transitions described respectively by a massless free Majorana fermion and a gauged Ising CFT. Notably, the Hamiltonian that realises the free Majorana fermion is analogous with the field theory Hamiltonian in addition to a density-density interaction.
The transition between bosonic Pfaffian and Halperin 220 [30]
The fuzzy sphere can be used to answer long-standing questions on fractional quantum Hall transitions. An example is the transition between the bosonic Moore-Read Pfaffian [105] and the Halperin 220 state [122] on a bilayer bosonic system, which has been predicted to be described by a massless free Majorana fermion with the fermion parity $\mathbb{Z}_{2}$ symmetry gauged [123], but lacks simulations. This work identifies the low-energy spectrum with the gauged Majorana fermion.
@@ 3.5 Studying Conformal Defects and Boundaries
Apart from the bulk CFTs, the fuzzy sphere can also be used to study their conformal defects and boundaries. Deforming a CFT with interactions living on a sub-dimensional defect may trigger a RG flow towards a non-trivial interacting IR fixed point. A defect IR theory with a smaller conformal symmetry is called a defect CFT (dCFT) [124, 125]. The dCFTs have rich physical structures, such as defect operators and bulk-to-defect correlation functions. Moreover, a bulk CFT can flow to several different dCFTs. Similarly, deformation on the boundary may trigger a flow towards a boundary CFT (bCFT). So far, the accessible defects/boundaries include the pinning-field defect of the 3D Ising CFT, in particular, its defect operator spectrum, correlators [6], $g$-function, defect-changing operators [9], its cusp [11], and the conformal boundaries of the 3D Ising CFT [12, 13]. Besides the defects and boundaries of 3D CFTs, one can also studying lower dimensional bulk CFTs on a (2+1)(2+1)D set-up, like the on a fuzzy thin torus [19].
Conformal pinning-field defect [6]
This is the first work that studies conformal defects with the fuzzy sphere. The simplest example of conformal defect is the pinning-field defect of the 3D Ising CFT [126, 127, 128, 129], where the defect line is completely polarised and the $\mathbb{Z}_{2}$ symmetry is explicitly broken. A defect line along $z$-direction that passes the origin point, after the radial quantisation, corresponds to the north and south poles of the sphere being polarised. Hence, to realise the pinning-field defect on the fuzzy sphere, one only needs to apply a pinning magnetic field to the north and south poles (Since only the $m=+s$ orbital has non-zero amplitude at the north pole and m=sm=-s at the south pole due to the locality, one only need to pin the $m=\pm s$ orbitals).
@@
This work studies the $g$-function of conformal defects and the defect-creation and changing operators. Similar to the central charge and the $F$-function in bulk CFTs, there exists a RG-monotonic quantity called the $g$-function for the line defects that is non-increasing along the flow [130, 131]. It is defined as the ratio between the partition functions of the defect CFT and the bulk CFT. On a different note, consider two semi-infinite pinning-field defects pinned towards opposite directions joint at one point, a defect-changing operator (also known as the domain wall operator) lives at the joining point [132, 133]. Similarly, a defect-creation operator lives at the end-point of a semi-infinite line defect. The relevance of the defect-changing operator is related to the stability of spontaneous symmetry breaking (SSB) on the line defect.
Zhou et al. realise the defect-creation and changing operators for the Ising pinning-field defect by acting a pinning field at the north pole, and opposite pinning fields at the north and south poles, respectively. The scaling dimensions are calculated through state-operator correspondence $\Delta_{\textrm{creation}}=0.108(5),\Delta_{\textrm{changing}}=0.84(5)$, indicating the instability of SSB on the Ising pinning-field defect. Moreover, they show that the $g$-function and many other CFT data can be calculated by taking the overlaps between the eigenstates of different defect configurations. Notably, this work has given the first non-perturbative result for the $g$-function $g=0.602(2)$.
@@
A cusp is two semi-infinite defect lines joined at one point at an angle. This can be realised on the fuzzy sphere through pinning fields at two points at an angle. Cuomo et al. study the cusps through various theoretical and numerical approaches. In particular, on the fuzzy sphere, they calculate the cusp anomalous dimension as a function of the angle for the Ising pinning-field defect, and verify its relation with the Zamolodchikov norm of the displacement operator.
@@
Besides the fuzzy sphere $S^{2}$, the regularisation with the lowest Landau level can also be used for other manifolds or dimensions. This work studies the 2D CFT on a “fuzzy circle.” Although Landau levels can only be defined on even space dimensions [137], one can reach odd space dimensions by compactifying one of the even dimensions. Specifically, Han et al. construct the LLL on a thin torus $T^{2}$ and sends one of the lengths to infinity while fixing the other. In this way, a circle $S^{1}$ is recovered in the thermodynamic limit. They construct the 2D Ising and 3-state Potts CFTs on the fuzzy circle and compare the operator spectrum and OPE coefficients with the Virasoro multiplet structure and the known results in the exactly solvable minimal models.